OpenStudy (memyselfandpi):

Please help me with more algebra one :) A quadratic equation is shown below: 3x2 − 16x + 2 = 0 Part A: Describe the solution(s) to the equation by just determining the radicand. Show your work. (5 points) Part B: Solve 9x2 + 3x − 2 = 0 by using an appropriate method. Show the steps of your work, and explain why you chose the method used.

OpenStudy (memyselfandpi):

Please help me :)

OpenStudy (stacey):

Do you know the quadratic formula?

jimthompson5910 (jim_thompson5910):

@MeMyselfAndPi hint: you'll use the discriminant formula \[\Large D = b^2 - 4ac\] for part A

OpenStudy (memyselfandpi):

Please help me :)

jimthompson5910 (jim_thompson5910):

did that hint help at all? are you able to determine the values of a,b,c?

OpenStudy (memyselfandpi):

Sorry for the delay..

OpenStudy (memyselfandpi):

Had to do something really fast..

OpenStudy (memyselfandpi):

Let me try to fill this in..

OpenStudy (memyselfandpi):

@jim_thompson5910 Sorry for late response, I filled it in and got (−16)2−4×3×2 which simplified to 232.

OpenStudy (memyselfandpi):

And doesn't that tell me that it has 2 solutions? (Since it is positive)

OpenStudy (stacey):

Yes.

OpenStudy (memyselfandpi):

Ok, on to Part B....

jimthompson5910 (jim_thompson5910):

yes, D = 232 so D > 0 meaning that we have 2 real solutions. The two solutions are different from each other. because D is not a perfect square, the solutions are going to be decimal numbers

OpenStudy (stacey):

Since 232 is not a perfect square, the solutions are irrational.

OpenStudy (memyselfandpi):

Alright.

OpenStudy (memyselfandpi):

So for Part B do I factor?

OpenStudy (memyselfandpi):

9x^2 + 3x − 2 = 0 (3x−1)(3x+2)=0 3x−1=0 or 3x+2=0

OpenStudy (memyselfandpi):

1/3 and -2/3?

OpenStudy (memyselfandpi):

I chose to factor because it seemed like the easiest way for me to do it...

OpenStudy (stacey):

Yes. You did that very well.

OpenStudy (memyselfandpi):

:)